# Surprise-Adaptive Intrinsic Motivation for Unsupervised Reinforcement Learning

### v1

@@ Abstract
Both entropy-minimizing and entropy-maximizing (curiosity) objectives for unsupervised reinforcement learning (RL) have been shown to be effective in different environments, depending on the environment’s level of natural entropy. However, neither method alone results in an agent that will consistently learn intelligent behavior across environments. In an effort to find a single entropy-based method that will encourage emergent behaviors in any environment, we propose an agent that can adapt its objective online, depending on the entropy conditions by framing the choice as a multi-armed bandit problem. We devise a novel intrinsic feedback signal for the bandit, which captures the agent’s ability to control the entropy in its environment. We demonstrate that such agents can learn to control entropy and exhibit emergent behaviors in both high- and low-entropy regimes and can learn skillful behaviors in benchmark tasks. Videos of the trained agents and summarized findings can be found on our project webpage.
@@ 1 Introduction
Unsupervised reinforcement learning (URL), or learning without access to an extrinsic reward function, has recently gained significant attention, often as a pretraining method (Jaderberg et al., 2017) or as a reward bonus in sparse reward domains (Schmidhuber, 1991; Pathak et al., 2017; Burda et al., 2019). A recent focus has been on developing objectives where the agent has no access to extrinsic rewards and instead develops emergent behaviors from an intrinsic motivation alone  (Lopes et al., 2012; Berseth et al., 2021; Kim et al., 2020). In this context, unsupervised RL holds the promise of being able to develop natural-like intelligence, i.e. generally-capable agents that can be deployed to solve diverse tasks across diverse environments. However, thus far, no single intrinsic motivation function has succeeded in capturing the complexity of motivation that gives rise to intelligent systems.
Interestingly, two seemingly opposing methods, surprise-minimization (Friston, 2010; Berseth et al., 2021) and surprise-maximization (curiosity) (Schmidhuber, 1991; Pathak et al., 2017), have been proposed as intrinsic motivations, with both methods performing well depending on the properties of the environment in which they are deployed. In general, surprise-minimizing methods (Berseth et al., 2021) perform well in environments with naturally high entropy that can be reduced through control, while curiosity-based methods (Pathak et al., 2017) are better suited to environments where explicit exploration is necessary to encounter novel information. However, both methods are known to possess failure modes when exposed to the opposite entropy regime (Schmidhuber, 2010; Sun & Firestone, 2020).
In this work, we propose an adaptive mechanism to select between maximizing and minimizing surprise in a given environment, based on the agent’s ability to exert control over its entropy conditions, which we frame as a multi-armed bandit problem. We experimentally validate our surprise-adaptive agent by demonstrating its ability to mirror a surprise-maximizing or -minimizing agent in didactic low- and high-entropy environments, respectively. We demonstrate more diverse emergent behaviors, as measured by the performance on extrinsic reward than observed from the single-objective agents.
@@ 2 Related work
There is a rich body of work in the field of unsupervised RL and intrinsic motivation, upon which our method builds. The most widely explored class of intrinsic objectives is related to improving state coverage through exploration bonuses. In low-dimensional settings, count-based methods (Auer, 2002; Bellemare et al., 2016; Machado et al., 2020) are simple and effective but do not always extend well to high-dimensions (Lobel et al., 2023). Another popular class of methods in high-dimensional settings uses prediction error as an exploration bonus (Schmidhuber, 1991; Pathak et al., 2017). Naive implementations of these novelty-seeking agents, however, can be susceptible to what is known as the "noisy-TV problem" (Schmidhuber, 2010), where the agent becomes transfixed by irreducible aleatoric noise in the environment. Various formulations have been developed to combat this issue, though issues often persist (Pathak et al., 2017; Houthooft et al., 2016; Burda et al., 2019).
An alternative class of intrinsic objectives incentivizes the agent to exert control over its environment. This class of methods is rooted in the free energy principle, a concept from neuroscience that posits that intelligent organisms seek out stable niches by learning to control their environment to minimize the entropy they experience over their lifetime (Friston, 2010). One prominent formulation in this class is that of empowerment, defined as the maximal mutual information between an agent’s actions and future states (Zhao et al., 2020; Karl et al., 2015; Klyubin et al., 2005). However, empowerment is computationally difficult to estimate in large or continuous state and action spaces. A more tractable approximation to the free-energy principle was proposed by Berseth et al. (2021) as surprise-minimization. In this formulation, an upper bound on an agent’s total trajectory entropy is minimized by rewarding the agent with the log-probability of the current state under the estimated state marginal distribution. This method has shown promising results in a diverse set of stochastic environments (Berseth et al., 2021; Rhinehart et al., 2021). Surprise-minimizing agents, however, can fall victim to the "dark room problem" (Sun & Firestone, 2020), where the agent discovers an area of the state-space without any stochastic dynamics and fails to seek out any additional experience.
@@
Prior works have explored adaptivity in RL and found that it can be beneficial for learning (Badia et al., 2020; Moskovitz et al., 2021). Similar to our work, Moskovitz et al. (2021) uses a multi-armed bandit to control a learning hyper-parameter. However, their method relies on extrinsic rewards for providing feedback to the bandit, while our method derives rewards based only on intrinsic signals.
@@ 3 Background
Entropy, surprise and curiosity
The notion of surprise derives from the optimization of the entropy of the state marginal distribution under the policy $\pi_{\phi}(a|s)$, which we denote $d^{\pi_{\phi}}(s_{t})$. Given an estimate of this state marginal distribution, $p_{\theta_{t-1}}(s_{t})$, we can express an estimate of the sum of the entropies of the state distribution across a trajectory (see Appendix A of (Berseth et al., 2021) for a full derivation):
@@
Recalling Equation 1, we can see that minimizing the sum of the state entropy over a trajectory (Equation 3) corresponds to an RL agent with a reward function given by:
@@
and maximizing this objective corresponds to an RL agent with a reward function given by:
@@
Traditional curiosity methods do not correspond directly to surprise-maximization as in Equation 5, however. Instead, these methods are generally defined in terms of prediction error in a forward dynamics model. For example, in Pathak et al. (2017), the intrinsic reward is defined as:
$r(s_{t},a_{t})=\frac{\eta}{2}||f(\phi(s_{t}),a_{t})-\phi(s_{t+1})||_{2}^{2}$ (6)
where $\phi(\cdot)$ is a learned representation of the state from an inverse dynamics model and $f$ is a learned forward dynamics model. These learned representations generally retain parameters across episodes, unlike our estimate of the state marginal distribution $p_{\theta_{t}}(s_{t+1})$.
However, we note that this formulation is conceptually similar to that expressed in Equation 5. Particularly, we note that if $p_{\theta_{t}}(s_{t+1})$ is a Gaussian distribution with mean $f(\phi(s_{t}),a_{t})=f(\phi(s_{t}))$ and $\phi$ is an identity mapping, then Equation 6 is equivalent to Equation 5.
@@ 4 Surprise-adaptive bandit
To design such an adaptive agent, we must first be able to optimize for either single-objective, which requires an estimation of the state marginal distribution at time $t$, parameterized by $\theta_{t}$ (denoted $p_{\theta_{t-1}}$ in Equation 3). In general, this estimation can be quite complex; In Berseth et al. (2021), the authors propose a simplification which we adopt here. The method estimates $\theta_{t}$ by first selecting an appropriate distribution family to model observations (i.e. Gaussian, Bernoulli, etc.) and using maximum likelihood estimation to estimate the sufficient statistics of the distributions, fitted to the history of observed states through time $t$. Further details on estimating the state marginal distribution are provided in Appendix A.
@@
To adaptively select between the two objectives online, we propose a multi-armed bandit approach. Provided with an appropriate performance signal, a bandit learns to select optimally between actions, trading off exploration with exploitation to minimize the overall regret, making it an appropriate choice for online adaption. The key question is what type of feedback is best to provide the bandit, given access only to intrinsic rewards. We propose a feedback mechanism grounded in the observation that the general goal in both surprise minimization and surprise maximization is for the agent to be able to affect a change in the level of surprise it experiences. In a low-entropy environment, the agent can best affect change by increasing entropy, and vice versa. Hence, the bandit should receive feedback that reflects this agency. We propose using the absolute percent difference between the entropy of the state marginal distribution at the end of the $m$th episode ($p_{\theta_{T}}^{(m)}$) and that of a random agent in the same environment ($p_{\theta_{\text{rand}}}$).
$f_{m}=\left|\frac{H(p^{(m)}_{\theta_{T}})-H(p_{\theta_{\text{rand}}})}{H(p_{\theta_{\text{rand}}})}\right|$ (7)
@@
The precise algorithm is as follows (Algorithm 1). At the start of training, we estimate the entropy of a random agent by collecting trajectories using a uniform random policy and averaging the entropy of the state marginal distributions, computed at the end of each trajectory (2). Then, at the start of each episode, we select an arm from the bandit according to the UCB algorithm (10), which determines if the agent will receive rewards according to Equation 4 or Equation 5 during the upcoming episode. The agent is trained for a single episode, using any RL algorithm (7). At the end of each episode, the bandit receives feedback $f_{m}$ on its selection (9). Algorithm 1 shows the full training procedure.
@@
1:Initialize network parameters $\phi$, replay buffer $\beta$, bandit parameters $\mu^{(0)}$, and $\alpha^{(0)}\sim\text{Bern}(0.5)$
2:Compute $H(p_{\theta_{\text{rand}}})$ by rolling out random trajectories
@@ 5 Experiments and analysis
To validate the usefulness and effectiveness of our method, we must demonstrate (1) Deficiencies in the single objective agents under particular entropy conditions and how these deficiencies arise from a lack of controllable entropy (2) Ability of the surprise-adaptive agent to select an objective based on the controllable entropy and to mimic the behavior of the single-objective agent (3) The ability of the surprise-adaptive agent to control entropy given high dimensional observations such as images, and finally (4) Correlation between entropy control and the emergence of intelligent behaviors.
With these goals in mind, we select several environments for evaluation; First, a set of didactic environments that are designed specifically to create low- and high-entropy conditions to demonstrate both the success and failure modes of single-objective entropy control. Second, a set of RL benchmark environments that are not selected with any particular entropy conditions in mind, and hence are demonstrative of how our algorithm could perform on arbitrary environments with unknown entropy conditions. For an environment with image-based observations we select Freeway from the original Atari games suite (bel, 2013)
For the high-entropy environments, we select the Tetris environment used in Berseth et al. (2021) and construct the new Butterflies environment, shown in Figure 1. In Tetris the agent must survive as long as possible by clearing rows of blocks before they reach the top of the frame. In Butterflies, the agent (red) must find and catch butterflies (blue) that are moving randomly in the map, within a fixed-length episode. For the low-entropy environment, we construct a static maze environment (Maze), in which the agent navigates around a map with a single goal state, for a fixed-length episode. For both Butterflies and Maze, we construct small (10x10) and large (32x32) versions of the environments. More details on these new environments are available in Section A.2. For the benchmark environments, we select the MinAtar (Young & Tian, 2019) suite of tasks. This suite consists of simplified versions of five Atari games, which are designed to make the state space fully observable.
Our analysis contrasts our method with the two dominant entropy-based intrinsic reward paradigms. Hence, we compare our method (S-Adapt) against an exclusively surprise-minimizing agent (S-Min) (Berseth et al., 2021) and an exclusively surprise-maximizing agent (S-Max). Here, the surprise-maximizing agent represents the space of curiosity methods (Pathak et al., 2017), though we note that our method could be implemented on top of any desired curiosity-based method. Additionally, as baselines, we compare the entropy-based intrinsic agents against an agent trained on the extrinsic reward (Extrinsic), and a random agent (Random).
@@ 5.1 Failures of Single-Objective Entropy Control
First, we demonstrate, qualitatively and quantitatively, the success and failure modes of single-objective entropy-based agents, using the didactic environments, indicating the need for an adaptive method.
Figure 1: The Butterflies (left) and Maze environments (right). S-Min trains the agent to actively catch the butterflies in order to prevent diverse state configurations while at the same time preventing the agent to navigate around Maze. S-Max trains the agent to avoid catching butterflies while navigating the Maze efficiently. These two didactic environments show that current intrinsic objectives fail to provide generally-useful objectives for RL agents and cannot adapt.
@@
Quantitatively, the S-Min agent achieves the lowest or near-lowest entropy in all environments, while the S-Max agent achieves the highest or near-highest entropy in all environments ( Figures 2, 3 and 4), as expected. However, we highlight that the qualitatively interesting direction for entropy control is correlated not with a single objective, but with the scale of the absolute difference in the final entropy achieved by the agent versus that of the Random agent (Figures 2, 3 and 4). In the Maze environment, the S-Max agent drives a significant increase in entropy over the Random agent, while the S-Min agent achieves a relatively small decrease (Figure 2). Similarly, in the Butterflies environment, the opposite holds in the large map (Figure 2(b)). Interestingly, in the small map, the S-Min and S-Max agents achieve roughly the same absolute change in entropy (Figure 2(a)). This is because in the smaller map, avoiding butterflies is equally challenging compared to catching butterflies, while in the larger map, the butterflies are easily avoided.
@@
Figure 3: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in the Butterflies environment. S-Min, Extrinsic and even the Random agent catch most of the butterflies in the small grid. Because of the small size of the grid, surprise-minimization and surprise-maximization are equally effective in entropy control, and hence the S-Adapt agent converges to S-Max. In the larger grid, however, the Random agent can’t catch many butterflies and hence has a high-entropy state distribution. Again, the S-Max learns to also avoid catching butterflies and the S-Min learns to catch butterflies. However, now the S-Min agent is able to cause a much more significant change in the state-marginal entropy. The S-Adapt agent identifies this and converges to S-Min, resulting in agents that catch more than half of the butterflies without access to the extrinsic reward.
Figure 4: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in Tetris. S-Min, S-Adapt, and the Extrinsic agent solve the game (i.e. consistently survive for more than 200 steps). Interestingly, the surprise-minimizing objective, which S-Adapt converges to, turns out to be a better learning signal than the row-clearing extrinsic reward in Tetris, as the learned policies are more stable and the average episodic surprise is the lowest.

### v2

@@ Abstract
Both entropy-minimizing and entropy-maximizing objectives for unsupervised reinforcement learning (RL) have been shown to be effective in different environments, depending on the environment’s level of natural entropy. However, neither method alone results in an agent that will consistently learn intelligent behavior across environments. In an effort to find a single entropy-based method that will encourage emergent behaviors in any environment, we propose an agent that can adapt its objective online, depending on the entropy conditions it faces in the environment, by framing the choice as a multi-armed bandit problem. We devise a novel intrinsic feedback signal for the bandit, which captures the agent’s ability to control the entropy in its environment. We demonstrate that such agents can learn to optimize task returns through entropy control alone in didactic environments for both high- and low-entropy regimes and learn skillful behaviors in certain benchmark tasks. Videos and summarized findings can be found on our project webpage [^1] Code is available at https://github.com/roger-creus/surprise-adaptive-agents.
@@ 1 Introduction
Unsupervised reinforcement learning (URL), or learning without access to an extrinsic reward function, has recently gained significant attention, often as a pretraining method (Jaderberg et al., 2017) or as a reward bonus in sparse reward domains (Schmidhuber, 1991; Pathak et al., 2017; Burda et al., 2019b). A recent focus has been on developing objectives where the agent has no access to extrinsic rewards and instead develops emergent behaviors from intrinsic motivation alone  (Lopes et al., 2012; Kim et al., 2020; Berseth et al., 2021). In this context, unsupervised RL holds the promise of being able to develop natural-like intelligence, i.e. generally-capable agents that can be deployed to solve diverse tasks across diverse environments. However, thus far, no single intrinsic motivation function has succeeded in capturing the complexity of motivation that gives rise to intelligent systems.
Interestingly, two seemingly opposing methods, surprise-minimization (Friston, 2010; Berseth et al., 2021) and surprise-maximization (Schmidhuber, 1991; Pathak et al., 2017; Hazan et al., 2019; Tiapkin et al., 2023), have been proposed as intrinsic motivations, with both methods performing well depending on the properties of the environment in which they are deployed. In general, surprise-minimizing methods (Berseth et al., 2021) perform well in environments with naturally high entropy that can be reduced through control, while curiosity-based methods (Pathak et al., 2017) are better suited to environments where explicit exploration is necessary to encounter novel information. However, both methods are known to possess failure modes when exposed to the opposite entropy regime (Schmidhuber, 2010; Sun & Firestone, 2020).
In this work, we propose an adaptive mechanism to select between maximizing and minimizing surprise in a given environment, based on the agent’s ability to exert control over its entropy conditions, which we frame as a multi-armed bandit problem. We experimentally validate our surprise-adaptive agent by demonstrating its ability to mirror a surprise-maximizing or -minimizing agent in didactic low- and high-entropy environments, respectively, and, in doing so, perform well on these tasks without any access to extrinsic task rewards. In benchmark environments, we demonstrate more diverse emergent behaviors, as measured by the performance on extrinsic task reward, than observed from the single-objective agents.
@@ 2 Related work
There is a rich body of work in the field of unsupervised RL and intrinsic motivation, upon which our method builds. The most widely explored class of intrinsic objectives is related to improving state coverage through exploration bonuses which reward some measure of novelty. In low-dimensional settings, count-based methods (Auer, 2002; Bellemare et al., 2016; Machado et al., 2020) are simple and effective but do not always extend well to higher dimensions (Lobel et al., 2023). Another popular class of methods in high-dimensional settings uses prediction error as an exploration bonus (Schmidhuber, 1991; Pathak et al., 2017). A conceptually similar idea is that of entropy maximization (Hazan et al., 2019; Tiapkin et al., 2023; Jain et al., 2023), which seeks to maximize the entropy of the distribution of states experienced by the agent throughout its lifetime. Naive implementations of these novelty-seeking agents, however, can be susceptible to what is known as the "noisy-TV problem" (Schmidhuber, 2010), where the agent becomes transfixed by irreducible aleatoric noise in the environment. Various formulations have been developed to combat this issue, though issues often persist (Houthooft et al., 2016; Pathak et al., 2017; Burda et al., 2019b). Though these methods are generally implemented as bonuses to the extrinsic reward, some works have also investigated the ability of curiosity-driven agents to achieve good task rewards without any access to the extrinsic reward (Burda et al., 2019a)
An alternative class of intrinsic objectives also targets the scenario where no extrinsic rewards are available by incentivizing the agent to exert control over its environment. This class of methods is rooted in the free energy principle, a concept from neuroscience that posits that intelligent organisms seek out stable niches by learning to control their environment to minimize the entropy they experience over their lifetime (Friston, 2010). One prominent formulation in this class is that of empowerment, defined as the maximal mutual information between an agent’s actions and future states (Klyubin et al., 2005; Karl et al., 2015; Zhao et al., 2020). However, empowerment is computationally difficult to estimate in large or continuous state and action spaces. A more tractable approximation to the free-energy principle was proposed by Berseth et al. (2021) as surprise-minimization. In this formulation, an upper bound on an agent’s total trajectory entropy is minimized by rewarding the agent with the log-probability of the current state under the estimated state marginal distribution. This method has shown promising results in a diverse set of stochastic environments (Berseth et al., 2021; Rhinehart et al., 2021). Surprise-minimizing agents, however, can fall victim to the "dark room problem" (Sun & Firestone, 2020), where the agent discovers an area of the state-space without any stochastic dynamics and fails to seek out any additional experience.
@@
Prior works have explored adaptivity in RL and found that it can be beneficial for learning (Badia et al., 2020; Moskovitz et al., 2021). Similar to our work, Moskovitz et al. (2021) uses a multi-armed bandit to control a learning hyperparameter. However, their method relies on extrinsic rewards for providing feedback to the bandit, while our method derives rewards based only on intrinsic signals.
@@ 3 Background
Entropy and surprise
The notion of surprise derives from the optimization of the entropy of the state marginal distribution under the policy $\pi_{\phi}(a|s)$, which we denote $d^{\pi_{\phi}}(s_{t})$. Given an estimate of this state marginal distribution, $p_{\theta_{t-1}}(s_{t})$, we can express an estimate of the sum of the entropies of the state distribution across a trajectory (see Appendix A of Berseth et al. (2021) for a full derivation):
@@
Recalling Equation 1, we can see that minimizing the sum of the state entropy over a trajectory (Equation 3) corresponds to a surprise minimizing agent (Berseth et al., 2021) with a reward function given by:
@@
Maximizing this objective corresponds to an RL agent with a reward function given by:
@@
which is similar to the rewards provided to the EntGame agent in Tiapkin et al. (2023).
@@ 4 Surprise-adaptive bandit
To design such an adaptive agent, we must first be able to optimize for either single-objective, which requires an estimation of the state marginal distribution at time $t$, parameterized by $\theta_{t}$ (denoted $p_{\theta_{t-1}}$ in Equation 3). In general, this estimation can be quite complex; In Berseth et al. (2021), the authors propose a simplification which we adopt here. The method estimates $\theta_{t}$ by first selecting an appropriate distribution family to model observations (i.e. Gaussian, Bernoulli, etc.) and using maximum likelihood estimation to estimate the sufficient statistics of the distributions, fitted to the history of observed states through time $t$. Further details on estimating the state marginal distribution as well as ablations on the choice of distribution are provided in Section A.3.
@@
To adaptively select between the two objectives online, we propose a multi-armed bandit approach. Provided with an appropriate performance signal, a bandit learns to select optimally between actions, trading off exploration with exploitation to minimize the overall regret, making it an appropriate choice for online adaption. The key question is what type of feedback is best to provide the bandit, given access only to intrinsic rewards. We propose a feedback mechanism grounded in the observation that the general goal in both surprise minimization and surprise maximization is for the agent to be able to affect a change in the level of surprise it experiences. In a low-entropy environment, the agent can best affect change by increasing entropy, and vice versa.
We propose using the absolute percent difference between the entropy of the state marginal distribution at the end of the $m$th episode ($H(p_{\theta_{T}}^{(m)})$) and that of a random agent in the same environment ($H(p_{\theta_{T}}^{\text{rand}})$) (Equation 6). The motivation for this is as follows: a random agent cannot control the environment entropy as it cannot take any actions in response to feedback. Agents that produce state visitation distributions whose entropy significantly diverges from that of a random agent must therefore be exerting control over the entropy in the environment. We therefore provide feedback to the bandit which promotes agents that can exert such control by rewarding large deviations from a random agent. Since we are approximating the state marginal distributions by an analytical distribution, we can compute $H(p_{\theta_{T}}^{(m)})$ analytically from the estimated parameters (see A.3 for further details).
$f_{m}=\left|\frac{H(p^{(m)}_{\theta_{T}})-H(p_{\theta_{T}}^{\text{rand}})}{H(p_{\theta_{T}}^{\text{rand}})}\right|$ (6)
@@
The precise algorithm is as follows (Algorithm 1). At the start of training, we estimate the entropy of a random agent by collecting trajectories using a uniform random policy and averaging the entropy of the state marginal distributions, computed at the end of each trajectory (2). Then, at the start of each episode $m$, we select an arm from the bandit, represented by binary indicator $\alpha^{(m)}$, according to the UCB algorithm (10), which determines if the agent will receive rewards according to Equation 4 or Equation 5 during the upcoming episode. The agent is trained for a single episode, using any RL algorithm (7). At the end of each episode, the bandit receives feedback $f_{m}$ on its selection (9). Algorithm 1 shows the full training procedure.
@@
1:Initialize network parameters $\phi$, replay buffer $\beta$, initial mean of bandit arms $\mu^{(0)}$, and initial optimization direction $\alpha^{(0)}\sim\text{Bern}(0.5)$
2:Compute $H(p_{\theta_{T}}^{{\text{rand}}})$ by rolling out random trajectories
@@ 5 Experiments and analysis
To validate the usefulness and effectiveness of our method, we must demonstrate (1) Deficiencies in the single objective agents under particular entropy conditions and how these deficiencies arise from a lack of controllable entropy (2) Ability of the surprise-adaptive agent to select an objective based on the controllable entropy and to mimic the behavior of the single-objective agents and, finally (3) Correlation between entropy control and the emergence of intelligent behaviors.
With these goals in mind, we select several environments for evaluation; First, a set of didactic environments that are designed specifically to create low- and high-entropy conditions to demonstrate both the success and failure modes of single-objective entropy control. Second, a set of RL benchmark environments that are not selected with any particular entropy conditions in mind, and hence are demonstrative of how our algorithm could perform on arbitrary environments with unknown entropy conditions.
For the high-entropy environments, we select the Tetris environment used in Berseth et al. (2021) and construct the new Butterflies environment, shown in Figure 1. In Tetris the agent must survive as long as possible by clearing rows of blocks before they reach the top of the frame. In Butterflies, the agent (red) must find and catch butterflies (blue) that are moving randomly in the map, within a fixed-length episode. For the low-entropy environment, we construct a static maze environment (Maze), in which the agent navigates around a map with a single goal state, for a fixed-length episode. For both Butterflies and Maze, we construct small (10x10) and large (32x32) versions of the environments. More details on these new environments are available in Section A.2. For the benchmark environments, we select the MinAtar (Young & Tian, 2019) suite of tasks. This suite consists of simplified versions of five Atari games, which are designed to make the state space categorical and fully observable without frame-stacking. Finally, to experiment on image-based observations, we test on Freeway from the original Atari games suite (Bellemare et al., 2013)
Our analysis contrasts our method with the two dominant entropy-based intrinsic reward paradigms. Hence, we compare our method (S-Adapt) against an exclusively surprise-minimizing agent (S-Min) (Berseth et al., 2021) and an exclusively surprise-maximizing agent (S-Max). Here, the surprise-maximizing agent represents the space of curiosity and maximum entropy methods (Pathak et al., 2017; Hazan et al., 2019), though we note that our method could be implemented on top of any desired curiosity-based method. Additionally, as baselines, we compare the entropy-based intrinsic agents against an agent trained on the extrinsic reward (Extrinsic), and a random agent (Random).
@@ 5.1 Failures of Single-Objective Entropy Control
First, we demonstrate, qualitatively and quantitatively, the success and failure modes of single-objective entropy-based agents, using the didactic environments.
@@
Figure 1: The Butterflies (left) and Maze environments (right). S-Min trains the agent to actively catch the butterflies in order to prevent diverse state configurations while at the same time preventing the agent to navigate around Maze. S-Max trains the agent to avoid catching butterflies while navigating the Maze efficiently. These two didactic environments show that current intrinsic objectives fail to provide generally useful objectives for RL agents and cannot adapt.
Quantitatively, we evaluate the average surprise and average extrinsic returns for the agents across training in all environments (Figures 2, 3 and 4). Notably, the S-Min agent achieves the lowest or near-lowest entropy in all environments, while the S-Max agent achieves the highest or near-highest entropy in all environments, as expected.
@@
However, we highlight that the qualitatively interesting direction for entropy control is correlated not with a single objective, but with the scale of the absolute difference in the final entropy achieved by the agent versus that of the Random agent. For example, in the Maze environment, the S-Max agent drives a significant increase in entropy over the Random agent, while the S-Min agent achieves a relatively small decrease (Figure 2). Similarly, in the Butterflies environment, the opposite holds in the large map (Figure 2(b)). Interestingly, in the small map, the S-Min and S-Max agents achieve roughly the same absolute change in entropy (Figure 2(a)). This is because in the smaller map, avoiding butterflies is equally challenging compared to catching butterflies, while in the larger map, the butterflies are easily avoided.
@@
Figure 3: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in the Butterflies environment. S-Min, Extrinsic and even the Random agent catch most of the butterflies in the small grid. Because of the small size of the grid, surprise-minimization and surprise-maximization are equally effective in entropy control, and hence the S-Adapt agent converges to S-Max. In the larger grid, however, the Random agent can’t catch many butterflies and hence has a high-entropy state distribution. Again, the S-Max agent learns to also avoid catching butterflies and the S-Min agent learns to catch butterflies. However, catching butterflies results in a significant change in the state-marginal entropy in this larger grid. The S-Adapt agent identifies this and converges to S-Min, resulting in agents that catch more than half of the butterflies without access to the extrinsic reward.
Figure 4: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in Tetris. S-Min, S-Adapt, and Extrinsic agents solve the game (i.e. consistently survive for more than 200 steps). Interestingly, the surprise-minimizing objective, which S-Adapt converges to, turns out to be a better learning signal than the row-clearing extrinsic reward in Tetris.